# Generalized Segal–Bargmann transform for Poisson distribution revisited

### v1

@@ 1 Introduction
A study of the Segal–Bargmann transform associated with a Gaussian measure on a (compex) infinite dimensional space was carried out by Y. M. Berezansky and Y. G. Kondratiev in their monograph [5, Chapter 2], see also the earlier work [12]. Essentially at the same time, the Segal–Bargmann transform in the infinite dimensional setting was discussed within white noise analysis under the name of $S$-transform. See e.g. the monographs [8, 17] and the references therein. In fact, in the infinite dimensional setting, it was natural to study the Segal–Bargmann transform of spaces of test and generalized functions.
@@
weakly converge to $\mu_{\sigma}$ as $\alpha\to 0$. Here $\mu_{\sigma}$ is the Gaussian distribution with mean zero and variance $\sigma$.
Let $(c_{n})_{n=0}^{\infty}$ be the monic polynomial sequence that is orthonal with respect to the measure $\mu_{\alpha,\sigma}$. In particular, for $\alpha=1$, $(c_{n})_{n=0}^{\infty}$ is a sequence of Charlier polynomials. The (exponential) generating function of the sequence $(c_{n})_{n=0}^{\infty}$ has the form
@@
We note that, in our recent preprint [11], we deal with the generalized Segal–Bargmann transform for the remaining Sheffer sequences of orthogonal polynomials in the classification of Meixner [16].
@@ 2.2 Sheffer operators
This section is based on [18, Chapter 3], [7] and [6].
@@ 3 Segal–Bargmann transform in the one-dimensional case
In this section, we follow [2] and [5, Chapter 2, Section 5.2] (in a slightly generalized form).
@@ 4 A generalized Segal–Bargmann transform for $\pi_{\alpha,\sigma}$
For $\alpha>0$ and $\sigma>0$, let the probability measure $\pi_{\alpha,\sigma}$ on $\alpha\mathbb{N}_{0}$ be defined by (1).
@@
$\lim_{\alpha\to 0}\int_{\alpha\mathbb{N}_{0}}e^{ixy}\,\tilde{\pi}_{\alpha,\sigma}(dy)=\exp\bigg{(}-\frac{\sigma}{2}\,y^{2}\bigg{)}=\int_{\mathbb{R}}e^{ixy}\,\mu_{\sigma}(dy).$
@@
Let $(c_{n})_{n=0}^{\infty}$ denote the monic polynomial sequence that is orthogonal with respect to $\pi_{\alpha,\sigma}$. The $(c_{n})_{n=0}^{\infty}$ is a Sheffer sequence having the generating function
The sequence $(c_{n})_{n=0}^{\infty}$ satisfies the recurrence formula
Formula (23) implies that $\|c_{n}\|_{L^{2}(\alpha\mathbb{N}_{0},\pi_{\alpha,\sigma})}^{2}=n!\,\sigma^{n}$. Hence, we can define a generalized Segal–Bargmann transform $\mathcal{S}:L^{2}(\alpha\mathbb{N}_{0},\pi_{\alpha,\sigma})\to\mathbb{F}_{\sigma}(\mathbb{C})$ as a unitary operator that satisfies $(\mathcal{S}c_{n})(z)=z^{n}$ for $n\in\mathbb{N}_{0}$.
@@
We use ideas similar to those in [11, Section 4]. We define linear operator $\mathcal{U}$ and $\mathcal{V}$ in $\mathbb{C}[z]$ by
@@
As easily seen, the operators $\mathcal{U}$, $\mathcal{V}$ satisfy the commutation relation $\mathcal{V},\mathcal{U}]=\alpha$. Hence, they are generators of a Weyl algebra, see e.g. [15, Section 5.6].
@@
In the case $\alpha=1$, the explicit form of the Charlier polynomials is well known. The reader is advised to compare formula (38) in the case $\alpha=1$ with an expansion of $\binom{z}{n}$ in the Charlier polynomials discussed in Example 1 on page 479 of [9]. Corollary 10 can also be derived from [11, Theorem 4.6] by taking the limit $\beta\to 0$.
@@ Acknowledgements
C.K. was financially supported by the EPSRC, UKRI as EPSRC Maths DTP Doctoral Prize related to grant reference EP/T517987/1 with finance code NCZ1071-100 which co-operated with Faculty of Science and Engineering, Swansea University.

### v2

@@ 1 Introduction
A study of the Segal–Bargmann transform associated with a Gaussian measure on a (complex) infinite dimensional space was carried out by Y. M. Berezansky and Y. G. Kondratiev in their monograph [6, Chapter 2], see also the earlier work [14]. Essentially at the same time, the Segal–Bargmann transform in the infinite dimensional setting was discussed within white noise analysis under the name of $S$-transform. See e.g. the monographs [10, 19] and the references therein. In fact, in the infinite dimensional setting, it was natural to study the Segal–Bargmann transform of spaces of test and generalized functions.
@@
weakly converge to $\mu_{\sigma}$ as $\alpha\to 0$. Here $\mu_{\sigma}$ is the Gaussian distribution with mean zero and variance $\sigma$. From the viewpoint of quantum physics, the interpolating parameter $\alpha$ provides a connection between the particle density of an infinite free Bose gas at zero temperature and a free Bose field, see Remark 7 below for detail.
Let $(c_{n})_{n=0}^{\infty}$ be the monic polynomial sequence that is orthogonal with respect to the measure $\mu_{\alpha,\sigma}$. In particular, for $\alpha=1$, $(c_{n})_{n=0}^{\infty}$ is a sequence of Charlier polynomials. The (exponential) generating function of the sequence $(c_{n})_{n=0}^{\infty}$ has the form
@@
We note that, in our recent paper [13], we dealt with the generalized Segal–Bargmann transform for the remaining Sheffer sequences of orthogonal polynomials in the classification of Meixner [18].
@@ 2.2 Sheffer operators
This section is based on [20, Chapter 3], [9] and [7].
@@ 3 Segal–Bargmann transform in the one-dimensional case
In this section, we follow [3] and [6, Chapter 2, Section 5.2] (in a slightly generalized form).
@@ 4 A generalized Segal–Bargmann transform for $\pi_{\alpha,\sigma}$
For $\alpha>0$ and $\sigma>0$, let the probability measure $\pi_{\alpha,\sigma}$ on $\alpha\mathbb{N}_{0}$ be defined by (1). Let $(c_{n})_{n=0}^{\infty}$ denote the monic polynomial sequence that is orthogonal with respect to $\pi_{\alpha,\sigma}$. The $(c_{n})_{n=0}^{\infty}$ is a Sheffer sequence having the generating function
The sequence $(c_{n})_{n=0}^{\infty}$ satisfies the recurrence formula
It follows from (23) that $\|c_{n}\|_{L^{2}(\alpha\mathbb{N}_{0},\pi_{\alpha,\sigma})}^{2}=n!\,\sigma^{n}$.
@@
$\lim_{\alpha\to 0}\int_{\alpha\mathbb{N}_{0}}e^{ixy}\,\tilde{\pi}_{\alpha,\sigma}(dy)=\exp\bigg(-\frac{\sigma}{2}\,y^{2}\bigg)=\int_{\mathbb{R}}e^{ixy}\,\mu_{\sigma}(dy).$ (24)
@@
Remark 7.
[…]
$\rho_{\alpha,\sigma}\,e_{n}=\sqrt{\sigma(n+1)}\,e_{n+1}+(\alpha n+\sigma/\alpha)e_{n}+\sqrt{\sigma n}\,e_{n-1},\quad n\in\mathbb{N}_{0}.$ (25)
Consider a creation operator $a^{+}$ and an annihilation operator $a^{-}$ in $\ell_{2}$ satisfying $a^{+}e_{n}=\sqrt{n+1}\,e_{n+1}$ and $a^{-}e_{n}=\sqrt{n}\,e_{n}$ ($n\in\mathbb{N}_{0}$). The operators $a^{+}$ and $a^{-}$ are adjoint of each other and satisfy the commutation relation $a^{-},a^{+}]=1$. By (25),
$\rho_{\alpha,\sigma}=\alpha(a^{+}+\sqrt{\sigma}/\alpha)(a^{-}+\sqrt{\sigma}/\alpha).$
Note that the operators $A_{\alpha,\sigma}^{+}=a^{+}+\sqrt{\sigma}/\alpha$ and $A_{\alpha,\sigma}^{-}=a^{-}+\sqrt{\sigma}/\alpha$ are also adjoint of each other and satisfy the commutation relation $A_{\alpha,\sigma}^{-},A_{\alpha,\sigma}^{+}]=1$. It follows from [1] (see also [8]) that the operators $A_{\alpha,\sigma}^{+}$, $A_{\alpha,\sigma}^{-}$ form a representation of the canonical commutation relations (CCR) describing an infinite free Bose gas at zero temperature. Hence, the operator $A_{\alpha,\sigma}^{+}A_{\alpha,\sigma}^{-}$ is the particle density of this gas with average density $\sigma/\alpha^{2}$. We observe that
$\rho_{\alpha,\sigma}-\frac{\sigma}{\alpha}$ $=\alpha\bigg(A_{\alpha,\sigma}^{+}A_{\alpha,\sigma}^{-}-\frac{\sigma}{\alpha^{2}}\bigg)$
$=\sqrt{\sigma}(a^{+}+a^{-})+\alpha a^{+}a^{-}\to\sqrt{\sigma}(a^{+}+a^{-})\quad\text{as }\alpha\to 0.$
The limiting operator $\sigma(a^{+}+a^{-})$ describes a free Bose field.
We define a generalized Segal–Bargmann transform $\mathcal{S}:L^{2}(\alpha\mathbb{N}_{0},\pi_{\alpha,\sigma})\to\mathbb{F}_{\sigma}(\mathbb{C})$ as a unitary operator that satisfies $(\mathcal{S}c_{n})(z)=z^{n}$ for $n\in\mathbb{N}_{0}$.
@@
We use ideas similar to those in [13, Section 4]. We define linear operator $\mathcal{U}$ and $\mathcal{V}$ in $\mathbb{C}[z]$ by
@@
As easily seen, the operators $\mathcal{U}$, $\mathcal{V}$ satisfy the commutation relation $\mathcal{V},\mathcal{U}]=\alpha$. Hence, they are generators of a Weyl algebra, see e.g. [17, Section 5.6].
@@
In the case $\alpha=1$, the explicit form of the Charlier polynomials is well known. The reader is advised to compare formula (40) in the case $\alpha=1$ with an expansion of $\binom{z}{n}$ in the Charlier polynomials discussed in Example 1 on page 479 of [11]. Corollary 11 can also be derived from [13, Theorem 4.6] by taking the limit $\beta\to 0$.
@@ Acknowledgements
The initial version of the paper was written when C.K. was a postdoc at Swansea University. The revision of the paper was conducted after C.K. had become a lecturer at Suranaree University of Technology.
C.K. was financially supported by the EPSRC, UKRI as EPSRC Maths DTP Doctoral Prize related to grant reference EP/T517987/1.
The authors are grateful to the anonymous refree for their helpful suggestion to improve the paper.